OpenStudy (anonymous):

help @privetek Find the value of x so that the line passing through (-2, 7) and (x, 2) has a slope of negative one third

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1) :) 1/3=(7-2)/(-2-x) 1/3(-2-x)=(7-2) -2/3 - 1/3x = 5 -1/3 x = 5 +2/3 x = (17/3)(-3) x = -17